Sterling Optical and Royal Optical both make glass frames and each is able to generate earnings before interest and taxes of $97,600. The separate capital structures for Sterling and Royal are shown here Sterling Royal Debt @ 8% Common stock, $5 par Total Common shares $. 732,000 Debt @ 8% $ 244,000 976,000 $1,220,000 195,200 488,000 Common stock, $5 par $1,220,000 Total 97,600 Common shares a. Compute earnings per share for both firms. Assume a 25 percent tax rate. (Round your answers to 2 decimal places.) Earnings per Share Sterling Royal b. In part a, you should have gotten the same answer for both companies earnings per share. Assuming a P/E ratio of 21 for each company, what would its stock price be? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Stock price c. Now as part of your analysis, assume the P/E ratio would be 15 for the riskier company in terms of heavy debt utilization in the capital structure and 24 for the less risky company. What would the stock prices for the two firms be under these assumptions? (Note Although interest rates also would likely be different based on risk, we will hold them constant for ease of analysis.) (Do not round intermediate calculations. Round your answers to 2 decimal places.) Stock Price Sterling Royal
Answer #1
Answer a
Computation of earnings per share for both firms
Sterling Royal
Earnings before interest and taxes $97,600.00 $97,600.00
Less : Interest Expense $58,560.00 $19,520.00
Earnings before taxes $39,040.00 $78,080.00
Less : Taxes @ 25% $9,760.00 $19,520.00
Net Income $29,280.00 $58,560.00
/ Number of common shares 97600 195200
Earnings Per share $0.30 $0.30
Answer b
Calculation of stock price
P/E ratio i.e.price earning ratio = Market price per share / Earnings per share
Hence , Stock price = P/E ratio * EPS
Stock price = 21 * $0.30 = $6.30
Answer c
Calculation of stock price of Sterling
Stock price = P/E ratio * EPS
Stock price = 15 * $0.30 = $4.50
Calculation of stock price of Royal
Stock price = P/E ratio * EPS
Stock price = 24 * $0.30 = $7.20